OpenStudy (anonymous):

What is a possible value for the missing term of the geometric sequence? 6, 30, 150, 750, ... a.148 b.4 c.41 d.2368

OpenStudy (amistre64):

how is a geometric sequence generated?

OpenStudy (amistre64):

i spose a random number generator could produce a geometric sequence ...

OpenStudy (amistre64):

seems to me like each value is getting larger then the last one ... so, the next one would have to be larger than 6, larger than 30, larger than 150, and larger than 750

OpenStudy (anonymous):

you multiply each time by 5

sammixboo (sammixboo):

The pattern for this is x5, so what is 750 x 5?

OpenStudy (amistre64):

its not asking for the next number, but logically out of the 4, which could it be

OpenStudy (anonymous):

it equals 3750

OpenStudy (anonymous):

i'd say d maybe

sammixboo (sammixboo):

Yes.

OpenStudy (anonymous):

is there a no answer available option?

OpenStudy (anonymous):

I know that. I just don't understand how to get the common ratio?? Is it 5?

OpenStudy (anonymous):

i am not sure

OpenStudy (amistre64):

well, since 6*r = 30; r = 30/6

OpenStudy (amistre64):

30*r = 150 .. r = 150/30

OpenStudy (anonymous):

Were finding the common ratio. We know that its geometric because it multiplies by 5

OpenStudy (cwrw238):

yea - i think theres a mistake here somewhere

OpenStudy (cwrw238):

none of these options can be inserted or added to make a geometric series